OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 5 is 3 more than –6. What is the number? A. –45 B. –30 C. –15 D. –5

pooja195 (pooja195):

Hi Welcome to OpenStudy :-) Can you start out by attempting to write the equation?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So

OpenStudy (anonymous):

I would take a negative number then divide it by 5 right?

pooja195 (pooja195):

No, first we need to write out equation.

OpenStudy (anonymous):

Ok So

OpenStudy (anonymous):

X Divided by 5= -3 right?

pooja195 (pooja195):

Where's the 6?

OpenStudy (anonymous):

It says 3 more then -6 so that would be -3

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ x }{ 5 }=-6+3\] Oh i see what you did now ok good :-) \[\huge~\rm~\bf~\frac{ x }{ 5 }=-3\] any ideas on what the next step is?

OpenStudy (anonymous):

Find out what X is

pooja195 (pooja195):

Good but what operation would we use?

OpenStudy (anonymous):

Um

pooja195 (pooja195):

We need to multiply by the reciprocal \[\huge~\rm~\bf~~\frac{ 5 }{ x }\times \frac{ x }{ 5 }=-3 \times 5 \] Like this.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

That would be -15 right?

pooja195 (pooja195):

Yes :) well done

OpenStudy (anonymous):

Thank you for the help

pooja195 (pooja195):

yw ^_^